3 2/3 (12-a)=43
I am a mom of a 7th grade student who is having difficulty with this problem. I come up with-2727273 as the answer but do not 1 know if it is correct, or how to explain how I got the answer.
Thank you
Judy Aucker

Answer by Alan3354(69443)

3 2/3 (12-a)=43
--------
If it's (3 2/3)*(12-a) = 43
Mixed numbers are useless.
(11/3)*(12-a) = 43
Multiply by 3 to eliminate the fractions.
11*(12-a) = 129
132 - 11a = 129
Subtract 132
-11a = -3
Divide by -11
a = 3/11